OpenStudy (anonymous):

Can someone help me with #32

OpenStudy (anonymous):

Here's the problem

jimthompson5910 (jim_thompson5910):

this is an isosceles right triangle so the two missing sides are both equal to x (unknown for now) 

jimthompson5910 (jim_thompson5910):

 we can use the Pythagorean theorem to tie all the sides of this right triangle together a^2 + b^2 = c^2 x^2 + x^2 = 24^2 2x^2 = 576 I'll let you finish. Tell me what you get when you solve for x